OpenStudy (anonymous):

this may be silly but how does 2*sqrt(2*g*h) = sqrt(2gh*2^2*h)

OpenStudy (anonymous):

2=sqrt(2^2) so 2*sqrt(2*g*h) =sqrt(2^2)*sqrt(2*g*h) =sqrt(2g*2^2*h)

OpenStudy (anonymous):

Thanks a bunch

OpenStudy (anonymous):

@fall2012 you cn close it now